OpenStudy (anonymous):

Cylinder A has a radius of 10 inches and a height of 5 inches. Cylinder B has a volume of 750π. What is the percent change in volume between cylinders A and B? (5 points) Select one: a. Cylinder B is 50% smaller than cylinder A. b. Cylinder B is 75% smaller than cylinder A. c. Cylinder B is 150% bigger than cylinder A. d. Cylinder B is 200% bigger than cylinder A.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

start by finding the Volume of cylinder A

ganeshie8 (ganeshie8):

do u have the formula handy ?

ganeshie8 (ganeshie8):

Volume of cylinder = \(\large \pi *r^2*h\)

OpenStudy (anonymous):

i forgot the formula all geometry is is formualas

OpenStudy (anonymous):

what is the radius and the height

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

ok i got it pi* 10^2*5

ganeshie8 (ganeshie8):

yes, so that equals 500pi

OpenStudy (anonymous):

so what am i supposed to do now

ganeshie8 (ganeshie8):

percent change = \(\large \frac{750\pi - 500\pi }{500 \pi} = \frac{250}{500} = \frac{1}{2}\)

ganeshie8 (ganeshie8):

multiply it wid 100 u wud get % change = 50

OpenStudy (anonymous):

so the answer is A

ganeshie8 (ganeshie8):

nope, A is not the answer

ganeshie8 (ganeshie8):

c. Cylinder B is 150% bigger than cylinder A.

ganeshie8 (ganeshie8):

100+50 = 150

OpenStudy (anonymous):

so C i totally get this:) thanks

ganeshie8 (ganeshie8):

good to hear :)